# Fundamental Oscillations of Massive Boson Stars - II

### v1

@@ 5.2 Comparison with other works
Static observables: Colpi et al. [12] first showed that in the strong-interaction limit, the EoS and, hence, the mass-radius curves are independent of the DM model parameters. Following this, many works have used this fact, but simple analytical expressions connecting mass, radius, and compactness were not available. Tang et al. [53] provided multiple fits for different mass regions for mini BS, i.e., in the Kaup limit. A relation between mass-compactness was reported in Pacilio et al. [56]. Using this and the $\Lambda-M$ relation reported in Sennett et al. [57], they also derived the $k_{2}-M$ relation. We update all these relations in this work.
Maselli et al. [41] showed that massive BSs follow the $I$-Love-$Q$ relations for select values of model parameters in the NS mass range. Wu et al. [87] also showed this for a few BS models. In this work, we explicitly prove that such universal relations are expected as the quantities do not scale with model parameters. We show this for $I$-Love-$C$ relations and do not consider the spin-induced quadrupole moment here as we restrict the non-spinning case, also providing a comparison with NSs.
Sennett et al. [57] reported that $\Lambda_{min}$ for BSs is 280. A recent work by Cipriani et al. [21] reported $\Lambda_{min}\approx 290$ and $k_{2,min}\approx 0.045$. Our results are consistent as we obtain $\Lambda_{min}=280$ and $k_{2,min}=0.045$. Cardoso et al. [88] studied tidal deformability in detail. The minimum value for the BSs considered here was not explicitly reported.
BS QNMs: [27] studied BS QNMs for odd parity, which have been shown not to couple to gravitational radiation. [28] explored the even parity QNM in detail, however, the work is restricted to non-interacting case ($\Lambda=0$). [29] explore QNMs of self-interacting BSs restricting themselves to weak-interaction limit ($\Lambda<200$) and do not focus on the damping times. [30] carry out a similar analysis but for three types of cases (Mini BS, Massive BS, and Solitonic BS). For the case of massive BS, as discussed in this work, they fix $\Lambda=200$, again in the weak interaction limit. The dimensionless quantity used was $\omega^{\prime}=\omega/m$ as opposed to $\omega^{\prime}=\omega\sqrt{\lambda}/m$ in our case. Thus, we cannot compare our results with those of these works. [32] studied radial QNMs, which do not give out GWs.
@@
Observations: On the observational front, [89, 90] provided an in-depth review of various observational signatures of ECOs using GW focussing on the inspiral phase of binary systems and detectability using LIGO was discussed. These were extended to include additional interaction terms [91] and for extreme mass ratio inspiral [92][34] extended this list to include quasinormal oscillations. Here, we discuss the detectability as well as parameter space and mass-compactness regions that can be probed using $f$-modes, in particular, that could also be excited in isolated BS using LIGO as well as future next-generation GW detectors.
@@ 5.3 Future Scope
Having derived the scaling and fit relations for non-radial $f$-mode oscillations, these can be used for the searches of massive BS on the strong-interaction limit where the effect of $f$-modes is important. These would be relevant in the future when $f$-mode oscillations become detectable. $f$-modes for BSs with different interaction potentials, like $\phi^{n}$ [93] or axion potentials, vector bosons, as well as other types of ECOs [5] need to be explored further in a similar way.
Stochastic GW background from binary systems of ECos [94], BSs [95], and other exotic sources [96] are also interesting prospects. The recent detection of stochastic GW by the IPTA has restored interest in stochastic GW sources. We found that we could have long-lived low-frequency $f$-modes for massive BSs depending on the DM model parameters. These BS oscillations could act as an additional source of the stochastic background. .
@@ Appendix A Scaling of Love equations
It is known that the Tolman-Oppenheimer-Volkoff (TOV) equations when written in terms of scaled parameters along the EoS, are independent of $m$ and $\lambda$ (see Appendix of [41]). The same holds for tidal deformability. This was shown in [57] that in the strong-interaction limit, the $\Lambda-M^{\prime}$ curves remain unchanged. We explicitly show and prove here that the solutions of electric and magnetic tidal love numbers are independent of model parameters when scaled dimensionless quantities are used.
@@
The tidal field consists of two parts [97] i) electric (even parity/polar) ($\epsilon_{L}$) and ii) magnetic (odd parity/axial) ($M_{L}$). The corresponding mass multipole ($Q_{L}$) and current multipole ($S_{L}$), which is purely GR effect, are related by the equations
$Q_{L}$ $=\lambda_{l}\epsilon_{L}~{},$ (A.1)
$S_{L}$ $=\sigma_{l}M_{L}~{},$ (A.2)
@@
where, $L$ are the space indices and $l$ is the order of multipole moment. Here, $\lambda_{l}$ and $\sigma_{l}$ are called the gravitoelectric and gravitomagnetic tidal deformability. Dimensionless tidal deformabilities ($\Lambda_{l},\Sigma_{l}$) are defined by dividing them by a factor $GM^{2l+1}$. These are given by
$\Lambda_{l}$ $=\frac{2}{(2l-1)!!}k_{l}\left(\frac{R}{GM}\right)^{2l+1}=\frac{2}{(2l-1)!!}\frac{k_{l}}{C^{2l+1}}~{},$ (A.3)
$\Sigma_{l}$ $=\frac{1}{4(2l-1)!!}j_{l}\left(\frac{R}{GM}\right)^{2l+1}=\frac{1}{4(2l-1)!!}\frac{j_{l}}{C^{2l+1}}$ (A.4)
@@
where $k_{l}$ and $j_{l}$ are the gravitoelectric and gravitomagnetic tidal Love numbers. These dimensionless tidal deformabilities depend on the EoS and on the mass configuration via the Love numbers and $C$. We want to see the effect of EoS on this, so we write the equations in terms of scaled quantities. We have already seen that when written in terms of scaled quantities $C=C^{\prime}$. The dimensionless tidal Love numbers are purely a function of dimensionless compactness and a dimensionless quantity y given by [97]
$k_{l}$ $=k_{l}(C,y_{l})~{},$ (A.6)
$j_{l}$ $=j_{l}(C,\tilde{y}_{l})~{}.$ (A.7)
which are given by
$y_{l}$ $=\frac{R}{H_{l}(R)}\frac{dH_{l}(R)}{dR}~{},$ (A.8)
$\tilde{y}_{l}$ $=\frac{R}{\tilde{H}_{l}(R)}\frac{d\tilde{H}_{l}(R)}{dR}$ (A.9)
Here, $R$ is the radius of the star. The functions $H_{l}$ and $\tilde{H}_{l}$ are obtained by solving the equations [97]
$\frac{d^{2}H_{l}(r)}{dr^{2}}$ $+\frac{dH_{l}(r)}{dr}\left(1-2\frac{Gm(r)}{r}\right)^{-1}\left[\frac{2}{r}-\frac{2Gm(r)}{r^{2}}-4\pi Gr(\rho(r)-p(r))\right]$
$+H_{l}(r)\left(1-2\frac{Gm(r)}{r}\right)^{-1}\Bigg{(}\Bigg{.}4\pi G\left[5\rho(r)+9p(r)+\dfrac{d\rho}{dp}(\rho(r)+p(r))\right]$
$-\frac{l(l+1)}{r^{2}}-4\left(1-\frac{2Gm(r)}{r}\right)^{-1}\left(\frac{Gm(r)}{r^{2}}+4\pi Grp(r)\right)^{2}\Bigg{.}\Bigg{)}=0,$ (A.10)
$\frac{d^{2}\tilde{H}_{l}(r)}{dr^{2}}$ $-\frac{d\tilde{H}_{l}(r)}{dr}\left(1-2\frac{Gm(r)}{r}\right)^{-1}4\pi Gr(\rho(r)+p(r))$
$-\tilde{H}_{l}(r)\left(1-2\frac{Gm(r)}{r}\right)^{-1}\Bigg{(}\Bigg{.}\frac{l(l+1)}{r^{2}}-\frac{4Gm(r)}{r^{3}}+8\pi G\theta(p(r)+\rho(r))\Bigg{.}\Bigg{)}=0~{}.$ (A.11)
Here, $c=1$ and $G=1/M_{pl}^{2}$. Here $\theta=+(-)1$ for purely static (irrotational) fluid. We now transform to scaled variable $r=r^{\prime}xM_{pl}$, $m=m^{\prime}xM_{Pl}^{3}$, $\rho=\rho^{\prime}/x^{2}$, and $p=p^{\prime}/x^{2}$ as discussed in Sec. 2. Then, these equations become
$\frac{d^{2}H_{l}(r^{\prime})}{dr^{\prime 2}x^{2}M_{Pl}^{2}}$ $+\frac{dH_{l}(r^{\prime})}{dr^{\prime}x^{2}M_{Pl}^{2}}\left(1-2\frac{m^{\prime}(r^{\prime})}{r^{\prime}}\right)^{-1}\left[\frac{2}{r^{\prime}}-\frac{2m^{\prime}(r^{\prime})}{r^{\prime 2}}-4\pi r^{\prime}(\rho^{\prime}(r^{\prime})-p^{\prime}(r^{\prime}))\right]$
$+\frac{H_{l}(r^{\prime})}{x^{2}M_{Pl}^{2}}\left(1-2\frac{m^{\prime}(r^{\prime})}{r^{\prime}}\right)^{-1}\Bigg{(}\Bigg{.}4\pi\left[5\rho^{\prime}(r^{\prime})+9p^{\prime}(r^{\prime})+\frac{d\rho^{\prime}}{dp^{\prime}}(\rho^{\prime}(r^{\prime})+p^{\prime}(r^{\prime}))\right]$
@@
$\frac{d^{2}\tilde{H^{\prime}}_{l}(r^{\prime})}{dr^{\prime 2}x^{2}M_{Pl}^{2}}$ $-\frac{d\tilde{H^{\prime}}_{l}(r^{\prime})}{dr^{\prime}x^{2}M_{Pl}^{2}}\left(1-2\frac{m^{\prime}(r^{\prime})}{r^{\prime}}\right)^{-1}4\pi r^{\prime}(\rho^{\prime}(r^{\prime})+p^{\prime}(r^{\prime}))$
$-\frac{\tilde{H^{\prime}}_{l}(r^{\prime})}{x^{2}M_{Pl}^{2}}\left(1-2\frac{m^{\prime}(r^{\prime})}{r^{\prime}}\right)^{-1}\Bigg{(}\Bigg{.}\frac{l(l+1)}{r^{\prime 2}}-\frac{4m^{\prime}(r^{\prime})}{r^{\prime 3}}+8\pi\theta(p^{\prime}(r^{\prime})+\rho^{\prime}(r^{\prime}))\Bigg{.}\Bigg{)}$ $=0~{}.$ (A.13)
@@
Thus, if we define $H^{\prime}_{l}=H_{l}/x^{2}M_{Pl}^{2}$ and $\tilde{H^{\prime}}_{l}=\tilde{H}_{l}/x^{2}M_{Pl}^{2}$, the equations become completely free of $x$, as we know from self-similar TOV that $m^{\prime}(r)$, $p^{\prime}(r^{\prime})$, $\epsilon^{\prime}(r^{\prime})$ are completely independent of $x$ and only depend on the central density. Thus, $H^{\prime}_{l}(r^{\prime})$ and $\tilde{H^{\prime}}(r^{\prime})$ are independent of $x$. This also means that $y^{\prime}_{l}(R^{\prime})$ and $\tilde{y^{\prime}}_{l}(R^{\prime})$ are independent of $x$. We can use these functions to calculate $y_{l}$ and $\tilde{y}_{l}$ for arbitrary EoS using
$y_{l}(R)$ $=\frac{R}{H_{l}(R)}\frac{dH_{l}(R)}{dR}=\frac{R^{\prime}}{H^{\prime}_{l}(R^{\prime})}\frac{dH^{\prime}_{l}(R^{\prime})}{dR^{\prime}}=y^{\prime}_{l}(R^{\prime}=R/xM_{Pl})~{},$ (A.14)
$\tilde{y}_{l}(R)$ $=\frac{R}{\tilde{H}_{l}(R)}\frac{d\tilde{H}_{l}(R)}{dR}=\frac{R^{\prime}}{\tilde{H^{\prime}}_{l}(R^{\prime})}\frac{d\tilde{H^{\prime}}_{l}(R^{\prime})}{dR^{\prime}}=\tilde{y^{\prime}}_{l}(R^{\prime}=R/xM_{Pl})$ (A.15)
@@
Thus, for a fixed $M^{\prime}/R^{\prime}/C$, the $y_{l}$, $\tilde{y}_{l}$ and consequently $k_{l}$, $j_{l}$, $\Lambda_{l}$ and $\Sigma_{l}$ are fixed and do no depend on $x$. $\Lambda_{l}(M^{\prime})$ and $\Sigma(M^{\prime})$ are thus EoS independent and can be used to obtain the corresponding quantities for arbitrary EoS using mass scaling. For example, the quadrupolar dimensionless tidal deformability of arbitrary EoS (parametrized by $x$) and mass configuration can be obtained as $\Lambda_{2}(x,M)=\Lambda_{2}(M^{\prime}=M/xM_{Pl}^{3})$. We have shown this for tidal Love numbers and dimensionless tidal deformability of all orders and both types.
Appendix B Variation with density
We perform another preliminary check in Fig. 14(a). We plot the mass ($M$), radius ($R$), and compactness ($C$) normalized to their respective maximum values reached for massive BSs as a function of scaled central density. Since the quantities are normalized, these curves are the same for BSs, as, for example, in the case of $M/M_{max}=M^{\prime}/M^{\prime}_{max}=M^{\prime}/0.06$. $R^{\prime}_{max}=0.626$ and $C_{max}=0.16$. Both mass and compactness increase with central density and asymptotically reach zero for low central densities. From the mass curve, we observe that beyond the critical central density $\rho^{\prime}_{c}=1.6$, $\partial M/\partial\rho_{c}$ becomes negative. Note that the maximum central density reached in a stable BS is $\mathcal{O}(1)$. Thus, the ultra-relativistic limit, where $\rho^{\prime}_{c}\gg 1,P^{\prime}=\rho^{\prime}/3$ is not reached. The compactness reaches a maximum beyond this density, but this is the region where BSs are unstable. Thus, the compactness at this critical density is the maximum compactness reached by stable BSs. We conclude that the maximum central density that can sustain stable configurations of BSs is given by $\rho_{c}=1.6/x^{2}$ for a given $(\lambda,m)$. Both mass and compactness undergo oscillations after having reached their maximum. Radius, on the other hand, starts from 1. This means that BSs reach a fixed maximum radius for low central densities. As the density increases, the radius reduces and reaches a minimum at a density beyond the critical density, which then oscillates. The radius at critical density is the minimum radius observed for a stable BS. This corresponds to $R^{\prime}=0.38$.
Figure 14: (a) The mass, radius, and compactness normalized by their maximum values as a function of dimensionless central energy density. The vertical dashed line denotes the maximum central density beyond which the star becomes unstable, i.e., $\partial M/\partial\rho_{c}<0$. This critical value is $\rho_{c}^{\prime}=1.6$. (b) Dimensionless tidal deformability as a function of scaled mass. The blue and green curves denote the first as provided in [56] and [55], respectively (c) $k_{2}$ as a function of scaled mass. The blue curve is derived from from $k_{2}=3/2\Lambda C^{5}$ using the fitting function C.1 as done in [56]. (d) Tidal love number $k_{2}$ as a function of C. The red curves are for a few NS EoSs as obtained from the CompOSE database [98]. The star in (b), (c), and (d) marks the point beyond which the BS configurations become unstable, as shown by the dashed curve.
Appendix C $\Lambda-M$
In Fig 14(b), we plot the dimensionless tidal deformability $\Lambda$ as a function of scaled mass. It has been shown the $\Lambda$ does not scale with $x$ in Appendix. A. Thus, we get a unique $\Lambda-M^{\prime}$ curve. $\Lambda$ decreases with mass, and we obtain a minimum value of $\Lambda_{min}=280$. We also show a fit given in Eq.9 of [56] This can be used to comment on the tidal deformability of BS with any model parameters. Cipriani et al. [21] mention in their work that for a given $x$, there is a range of masses for which the tidal deformability is measurable. This $\Lambda-M$ relation, along with scaling relations, can be used to get an analytical expression for the same. Vaglio et al. [55] also reported a fit in the form $\Lambda=\Lambda(M^{\prime})$ (see. Eqs. 11-12 from [55]). These fits are shown in Fig 14(b) for comparison.
For the $l=2$ tidal love number $k_{2}$, we obtain a minimum value of 0.045 for the highest compactness configuration, and $k_{2}$ approaches 0.25 for low compactness. This range is consistent with [56]. For low compactness, the central density for BSs is $\rho^{\prime}_{c}\ll 1\implies P^{\prime}=\rho^{\prime 2}/4$. Thus, it follows the polytrope relation $P=K\rho^{2}$. This is consistent with $k_{2}$ for a polytropic EoS $P=K\rho^{2}$ evaluated in the Newtonian limit, given by $k_{2}=\frac{15-\pi^{2}}{2\pi^{2}}\approx 0.26$ [99]. We plot $k_{2}$ as a function of scaled mass in Fig 14(c) We point out that our result does not match with $k_{2}$ evaluated in [56] (shown by blue curve in Fig 14(c)) as they used the fitting function between $C$ and $M$ given by
$C^{-1}\approx 7.5+48.8\left(1-\frac{M^{\prime}}{0.06}\right)^{2}$ (C.1)
to evaluate $k_{2}$ as a function of mass from the relation $k_{2}=\Lambda C^{5}$. As we showed in Fig. 2(a), this $C(M)$ relation is not a good fit for BSs. Moreover, a factor of 3/2 was missing in the expression for $k_{2}$ ($k_{2}=\frac{3}{2}\Lambda C^{5}$), which we have accounted for in the comparison. This is why Fig. 6 of [56] is inconsistent with our results and that from [21]. Also, the tidal love number $k_{2}$ as a function of compactness is shown in Fig. 14(d). We show the relation for NSs in red using a few NS EoSs taken from CompOSE [98] for comparison.
Appendix D Scaling for Moment of Inertia
The moment of inertia is defined as [100]
$I\equiv\frac{S}{\Omega}~{},$ (D.1)

### v3

@@ 5.2 Comparison with other works
Static observables: Colpi et al. [Colpi1986] first showed that in the strong-interaction limit, the EoS and, hence, the mass-radius curves are independent of the DM model parameters. Following this, many works have used this fact, but simple analytical expressions connecting mass, radius, and compactness were not provided. Tang et al. [Yang2024] provided multiple fits for different mass regions for mini BSs, i.e., in the Kaup limit. Pacilio et al. [Pacilio2020] reported an analytical relation between mass-compactness. None of these works derived analytical fits for macroscopic properties of massive BSs, as provided in this work.
Maselli et al. [Maselli2017] showed that massive BSs follow the II-Love-$Q$ relations for select values of model parameters in the NS mass range. Wu et al. [Wu2023] also showed this for a few BS models. In this work, we explicitly prove that such universal relations are expected as the quantities do not scale with model parameters. We show this for II-Love-$C$ relations and do not consider the spin-induced quadrupole moment here as we restrict the non-spinning case, also providing a comparison with NSs.
Sennett et al. [Sennett2017] reported that $\bar{\Lambda}_{\rm min}$ for BSs is 280. A recent work by Cipriani et al. [Cipriani2024] reported $\bar{\Lambda}_{\rm min}\approx 290$ and $k_{2,\rm min}\approx 0.045$. Our results are consistent as we obtain $\bar{\Lambda}_{\rm min}=280$ and $k_{2,\rm min}=0.045$. Cardoso et al. [Cardoso2017] studied tidal deformability in detail. The minimum value for the BSs considered here was not explicitly reported.
BS QNMs: There are several studies of BS QNMs in the literature with which a comparison cannot be made. We mention these studies here for completeness, along with the reason why the comparison is infeasible. [Kojima1991] studied BS QNMs for odd parity, which have been shown not to couple to gravitational radiation. [Yoshida1994] explored the even parity QNMs in detail; however, the work is restricted to the non-interacting case ($\Lambda=0$). [Balakrishna1998] explores QNMs of self-interacting BSs, restricting themselves to the weak-interaction limit ($\Lambda<200$), and does not focus on the damping times. [Macedo2013a] carries out a similar analysis but for three types of cases (Mini BS, Massive BS, and Solitonic BS). For the case of massive BSs, as discussed in this work, they fix $\Lambda=200$, again in the weak interaction limit. The dimensionless quantity used was $\omega^{\prime}=\omega/m$ as opposed to $\omega^{\prime}=\omega\sqrt{\lambda}/m$ in our case. Thus, we cannot compare our results with those of these works. [Kain2021] studied radial QNMs, which do not give out GWs.
@@
Observations: On the observational front, [Guidice2016, Bezares2018] provided an in-depth review of various observational signatures of ECOs using GWs, focussing on the inspiral phase of binary systems, and the detectability using LIGO was discussed. These were extended to include additional interaction terms [Croon2019] and for extreme mass ratio inspiral [Guo2019][shirke_BS_fmodes_short] extended this list to include quasinormal oscillations. Here, we discuss the detectability as well as parameter space and mass-compactness regions that can be probed using BS $f$-modes using LIGO as well as future next-generation GW detectors.
@@ 5.3 Future Scope
Having derived the scaling and fit relations for non-radial $f$-mode oscillations, these can be used for the searches of massive BSs in the strong-interaction limit where the effect of $f$-modes is important. These would be relevant in the future when $f$-mode oscillations become detectable. $f$-modes for BSs with different interaction potentials, like $\phi^{n}$ [Pitz2023] or axion potentials, vector bosons, as well as other types of ECOs [CardosoPani2019] need to be explored further in a similar way. Our analysis on detectability also implies that the non-detection of such bursts can, in turn, be used to rule out BS bursts with parameters indicated in Figs. 8 and 9.
Stochastic GW background from binary systems of ECOs [Barausse2018], BSs [Croon2018], and other exotic sources [Banks2023] are also interesting prospects. The recent detection of stochastic GWs by the IPTA has restored interest in stochastic GW sources. We found that we could have long-lived low-frequency $f$-modes for massive BSs depending on the DM model parameters. These BS oscillations could act as an additional source of the stochastic background.
@@ Appendix A Scaling of Love equations
It is known that the Tolman-Oppenheimer-Volkoff (TOV) equations, when written in terms of scaled parameters, are independent of $m$ and $\lambda$ (see Appendix of [Maselli2017]). The same holds for tidal deformability. This was shown in [Sennett2017] that in the strong-interaction limit, the $\bar{\Lambda}-M^{\prime}$ curves remain unchanged. We explicitly show and prove here that the solutions of electric and magnetic tidal Love numbers are independent of model parameters when scaled dimensionless quantities are used. We show this by considering the usual definition of Love numbers and the formalism used to compute them for any compact object given the TOV profile $\rho(r)$, $p(r)$, $m(r)$ as shown below. Note that the scaling is valid only in the strong interaction limit.
@@
The tidal field consists of two parts [Perot2021] i) electric (even parity/polar) ($\epsilon_{L}$) and ii) magnetic (odd parity/axial) ($M_{L}$). The corresponding mass multipole ($Q_{L}$) and current multipole ($S_{L}$), a pure GR effect, are related by the equations
$Q_{L}$ $=\bar{\lambda_{l}}\epsilon_{L}~,~S_{L}=\bar{\sigma_{l}}M_{L}~,$ (A.1)
@@
where, $L$ are the space indices and $l$ is the order of multipole moment. Here, $\bar{\lambda_{l}}$ and $\bar{\sigma_{l}}$ are called the gravitoelectric and gravitomagnetic tidal deformability. Dimensionless tidal deformabilities ($\bar{\Lambda_{l}},\bar{\Sigma_{l}}$) are defined by dividing them by a factor $GM^{2l+1}$. These are given by
$\bar{\Lambda_{l}}$ $=\frac{2}{(2l-1)!!}k_{l}\left(\frac{R}{GM}\right)^{2l+1}=\frac{2}{(2l-1)!!}\frac{k_{l}}{C^{2l+1}}~,$ (A.2)
$\bar{\Sigma_{l}}$ $=\frac{1}{4(2l-1)!!}j_{l}\left(\frac{R}{GM}\right)^{2l+1}=\frac{1}{4(2l-1)!!}\frac{j_{l}}{C^{2l+1}}$ (A.3)
@@
where $k_{l}$ and $j_{l}$ are the gravitoelectric and gravitomagnetic tidal Love numbers. These dimensionless tidal deformabilities depend on the EoS and on the mass configuration via the Love numbers and $C$. We want to see the effect of EoS on this, so we write the equations in terms of scaled quantities. We have already seen that when written in terms of scaled quantities $C=C^{\prime}$. The dimensionless tidal Love numbers are purely a function of dimensionless compactness and a dimensionless quantity $y$ given by [Perot2021] $k_{l}=k_{l}(C,y_{l});~j_{l}=j_{l}(C,\tilde{y}_{l})~.$ $y_{l}$ and $\tilde{y}_{l}$ are in turned evaluated as $y_{l}=\frac{R}{H_{l}(R)}\frac{dH_{l}(R)}{dR};~\tilde{y}_{l}=\frac{R}{\tilde{H}_{l}(R)}\frac{d\tilde{H}_{l}(R)}{dR}$, where the profiles for $H_{l}$ and $\tilde{H}_{l}$ are computed by solving the differential Eqs. 10 and 11 from [Perot2021]. Here, $R$ is the radius of the star. By applying the scaling relations $r=r^{\prime}xM_{\rm Pl}$, $m=m^{\prime}xM_{\rm Pl}^{3}$, $\rho=\rho^{\prime}/x^{2}$, and $p=p^{\prime}/x^{2}$ as discussed in Sec. 2 for massive BSs in the strong-interaction limit, these equations become
$\frac{d^{2}H_{l}(r^{\prime})}{dr^{\prime 2}x^{2}M_{\rm Pl}^{2}}$ $+\frac{dH_{l}(r^{\prime})}{dr^{\prime}x^{2}M_{\rm Pl}^{2}}\left(1-2\frac{m^{\prime}(r^{\prime})}{r^{\prime}}\right)^{-1}\left[\frac{2}{r^{\prime}}-\frac{2m^{\prime}(r^{\prime})}{r^{\prime 2}}-4\pi r^{\prime}(\rho^{\prime}(r^{\prime})-p^{\prime}(r^{\prime}))\right]$
$+\frac{H_{l}(r^{\prime})}{x^{2}M_{\rm Pl}^{2}}\left(1-2\frac{m^{\prime}(r^{\prime})}{r^{\prime}}\right)^{-1}\Bigg(\Bigg.4\pi\left[5\rho^{\prime}(r^{\prime})+9p^{\prime}(r^{\prime})+\frac{d\rho^{\prime}}{dp^{\prime}}(\rho^{\prime}(r^{\prime})+p^{\prime}(r^{\prime}))\right]$
@@
$\frac{d^{2}\tilde{H}_{l}(r^{\prime})}{dr^{\prime 2}x^{2}M_{\rm Pl}^{2}}$ $-\frac{d\tilde{H}_{l}(r^{\prime})}{dr^{\prime}x^{2}M_{\rm Pl}^{2}}\left(1-2\frac{m^{\prime}(r^{\prime})}{r^{\prime}}\right)^{-1}4\pi r^{\prime}(\rho^{\prime}(r^{\prime})+p^{\prime}(r^{\prime}))$
[…]
@@
Thus, if we define $H^{\prime}_{l}=H_{l}/x^{2}M_{\rm Pl}^{2}$ and $\tilde{H^{\prime}}_{l}=\tilde{H}_{l}/x^{2}M_{\rm Pl}^{2}$, the equations become completely free of $x$, as we know from self-similar TOV that $m^{\prime}(r)$, $p^{\prime}(r^{\prime})$, $\epsilon^{\prime}(r^{\prime})$ are completely independent of $x$ and only depend on the central density, i.e, we obtain the same Eqs. 10 and 11 of [Perot2021] in term of $H^{\prime}_{l}$ and $\tilde{H^{\prime}}_{l}$.
@@
Thus, profiles $H^{\prime}_{l}(r^{\prime})$ and $\tilde{H^{\prime}}(r^{\prime})$ are independent of $x$. This also means that $y^{\prime}_{l}(R^{\prime})$ and $\tilde{y}^{\prime}_{l}(R^{\prime})$ are independent of $x$. Thus, like $C$, we get a single value for both $y^{\prime}_{l}(R^{\prime})$ and $\tilde{y}^{\prime}_{l}(R^{\prime})$, independent of $x$ for a given central density $\rho^{\prime}$. This translates to single value of $k_{l}$, $j_{l}$ and $\bar{\Lambda}_{l}$, $\bar{\Sigma}_{l}$. We can then vary the central density to get a unique relation between the dimensionless tidal deformabilities and $C$ or $M^{\prime}$. In other words, for a fixed $M^{\prime}/R^{\prime}/C$, the $y_{l}$, $\tilde{y}_{l}$ and consequently $k_{l}$, $j_{l}$, $\bar{\Lambda_{l}}$ and $\bar{\Sigma_{l}}$ are fixed and do no depend on $x$. $\bar{\Lambda_{l}}(M^{\prime})$ and $\bar{\Sigma}(M^{\prime})$ are thus EoS independent and can be used to obtain the corresponding $\bar{\Lambda_{l}}(M)$ and $\bar{\Sigma}(M)$ for arbitrary EoS using mass scaling. For example, the quadrupolar dimensionless tidal deformability of arbitrary EoS (parametrized by $x$) and mass configuration can be obtained as $\bar{\Lambda_{2}}(x,M)=\bar{\Lambda_{2}}(M^{\prime}=M/xM_{\rm Pl}^{3})$. We have denoted $\bar{\Lambda_{2}}$ by $\bar{\Lambda}$ throughout the paper. We have mathematically shown that the tidal Love numbers and dimensionless tidal deformability of all orders and both types do not scale with $x$.
Appendix B Scaling for Moment of Inertia
The moment of inertia is defined as [YagiYunes2013]
$I\equiv\frac{S}{\Omega}~,$ (B.1)